# Attention Normalization Impacts Cardinality Generalization in Slot Attention

### v1

@@ Abstract
Object-centric scene decompositions are important representations for downstream tasks in fields such as computer vision and robotics. The recently proposed Slot Attention module, already leveraged by several derivative works for image segmentation and object tracking in videos, is a deep learning component which performs unsupervised object-centric scene decomposition on input images. It is based on an attention architecture, in which latent slot vectors, which hold compressed information on objects, attend to localized perceptual features from the input image. In this paper, we show that design decisions on normalizing the aggregated values in the attention architecture have considerable impact on the capabilities of Slot Attention to generalize to a higher number of slots and objects as seen during training. We argue that the original Slot Attention normalization scheme discards information on the prior assignment probability of pixels to slots, which impairs its generalization capabilities. Based on these findings, we propose and investigate alternative normalization approaches which increase the generalization capabilities of Slot Attention to varying slot and object counts, resulting in performance gains on the task of unsupervised image segmentation.
© 2024, the authors. All rights reserved.
@@ 2.1 Slot Attention
For the attention mechanism, an unnormalized $N\times K$ matrix $\bm{M}$ of dot products is formed from the keys $\bm{k}_{n}:=k(\bm{x}_{n})$ and queries $\bm{q}_{k}:=q(\bm{\theta}_{k})$. On each row of $\bm{M}$, a Softmax operator is then applied, yielding $\bm{\Gamma}=(\gamma_{n,k})\in[0,1]^{N\times K}$: $M_{n,k}:=\frac{1}{\tau}k(\bm{x}_{n})^{\top}q(\bm{\theta}_{k})=\frac{1}{\tau}\bm{k}_{n}^{\top}\bm{q}_{k}$ (1) $\gamma_{n,k}:=\frac{\exp{M_{n,k}}}{\sum_{k^{\prime}=1}^{K}\exp{M_{n,k^{\prime}}}}.$ (2) With this, each row $\bm{\gamma}_{n,:}$ may be interpreted as the probability of an input token $n$ to be assigned to a particular slot $k$. The constant $\tau$ corresponds to a temperature parameter which is chosen to be $\sqrt{D}$. A linear map $v:\mathbb{R}^{D_{\text{input}}}\to\mathbb{R}^{D}$ extracts values from the input tokens and the matrix $\bm{\Gamma}$ is used to accumulate values into unnormalized slot-wise update codes:
@@ 2.2 Von Mises-Fisher Distributions
Von Mises-Fisher (vMF) distributions (Fisher, 1953) are probability distributions on the unit $(d-1)$-sphere in $\mathbb{R}^{d}$. Typically, they are parametrized by a mean direction $\bm{\theta}\in\mathbb{R}^{d}$ with $\|\bm{\theta}\|_{2}=1$ and a concentration parameter $\tau>0$. They are defined by the following density w.r.t. the usual surface measure on the $(d-1)$-sphere:
$f(\bm{x}\;|\;\bm{\theta},\tau)=\frac{1}{Z(d,\tau)}\exp\left(\frac{\bm{\theta}^{\top}\bm{x}}{\tau}\right)$ (5)
where $Z(d,\tau)$ is a normalization constant that is independent of $\bm{\theta}$. If $(\bm{\theta}_{1},...,\bm{\theta}_{K})$ and $(\tau_{1},...,\tau_{K})$ are parameters of vMF distributions and $(\pi_{1},\dots,\pi_{K})$ is contained in the probability simplex, a vMF mixture model can be defined as usual via the density:
@@ 3 Slot Attention and von Mises-Fisher Mixture Model Parameter Estimation
Many works (Locatello et al., 2020; Chang et al., 2022b; a; Kirilenko et al., 2023) compare Slot Attention to expectation maximization (Dempster et al., 1977; Bishop, 2006) (EM) in Gaussian mixture models, i.e. to soft k-means clustering. We, however, connect it with expectation maximization in a mixture model of von Mises-Fisher (vMF) distributions, since Slot Attention uses a bilinear form on slots and inputs as a scoring function instead of the negative Euclidean distance. In this section, we make the parallel between Slot Attention and EM explicit by performing EM parameter estimation in a vMF mixture model and relating each step to the corresponding step in Slot Attention. We will then view the weighted mean, layer norm, and weighted sum normalization variants in the context of this analogy and compare them.
@@ 3.2 Comparing Normalizations in EM Analogy
In the weighted mean case, the aggregated values $\bm{u}_{k}$ can hold sufficient information to extract the quantities in (9) and (10) if $D={D_{\text{input}}}$ holds. Assuming that the value map is the identity, the right-hand side of (9) may be computed as $\bm{u}_{k}/\|\bm{u}_{k}\|_{2}$. However, it is not clear how $\pi_{k}$ could be computed from $\bm{u}_{k}$. Indeed, we show in Proposition 1 that there can be no general formula as for the weighted sum case that generalizes without exception across slot-counts. We provide a proof in Appendix C.
@@
for arbitrary $K\geq 1$, arbitrary slots $\tilde{\bm{\theta}}_{1},...,\tilde{\bm{\theta}}_{K}$ and resulting normalized update codes $\bm{u}_{1},...,\bm{u}_{k}$. I.e., there exists no function that infers column sums of $\bm{\Gamma}$ from the corresponding normalized update codes.
@@
Since the assumptions of Proposition 2 only exclude a parameter subset of zero volume, its conclusion likely holds in practice during training. Hence, the weighted sum variant may also be seen as a generalization of the weighted mean normalization: the update codes from weighted sum normalization hold sufficient information such that weighted-mean-normalized update codes can be recovered from them (e.g. by the update network if it has sufficient capacity). The reverse is not true, as demonstrated in Proposition 1.
@@ 4.1 Weighted Sum Normalization with Fixed Scaling
Here, we first use the triangle inequality, followed by the crude estimate $|v_{n,d}|\leq\max_{n}|v_{n,d}|$ for all $n$. Finally, we used the fact that $\sum_{n=1}^{N}\gamma_{n,k}\leq N$ holds, since $\bm{\Gamma}$ is row-stochastic with $N$ rows
@@ 4.2 Weighted Sum Normalization with Batch Scaling
where $\epsilon>0$ is a small constant. We stress that the values $m$ and $v$ are computed for each mini-batch in the first Slot Attention iteration and are therefore independent of $j$. Moreover, gradients flow through $m$ and $v$. We cache an exponential moving average of the batch statistics during training and use it during inference. Hence, during inference, the normalization in equation (15) is simply an affine transformation with fixed weights, thereby closely resembling the weighted sum normalization presented in the previous subsection. While the weighted sum normalization is linear and not affine, we note that this distinction does not impact the capacity of the model. Indeed, the vectors $\bm{u}_{k}$ are also affinely transformed within the update network, therefore making any previous affine or linear transformation redundant from a capacity perspective. Notwithstanding, the normalizations presented here will significantly alter the trainig trajectory of the models. Batch normalization, in particular, has previously been shown to remedy the problem of saturating activations and vanishing gradients (Ioffe & Szegedy, 2015; Pascanu et al., 2013), which may arise from improper normalization (Glorot & Bengio, 2010).
@@ 5 Experiments
We investigate the proposed normalizations on unsupervised object discovery tasks. To this end, we train autoencoders on the CLEVR (Johnson et al., 2017) and MOVi-C (Greff et al., 2022) datasets, utilizing autoencoder architectures that have been described in (Locatello et al., 2020) and (Seitzer et al., 2023), respectively. We will give a brief overview on our experimental setup in the following and refer to the supplementary material for more details.
@@ 5.1 Object Discovery on CLEVR
In this first set of experiments, we follow (Locatello et al., 2020) as closely as possible and train the previously described CNN-based architecture on the CLEVR6 dataset with 7 slots. We compare the baseline and layer normalizations proposed in (Locatello et al., 2020) to the methods discussed in Section 4. For each variant, we perform 5 training runs with different seeds. The trained models are evaluated on the CLEVR6 and CLEVR10 test sets. The baseline and layer norm variants lead to object-centric behavior for all 5 seeds. For the weighted sum normalization and the batch norm variant, however, we encounter two runs each in which the autoencoders decompose the input spatially instead of object-wise. Following (Locatello et al., 2020), we omit these runs in our analysis.
@@
In Subfigures 1(a) and 1(b), we illustrate how the foreground segmentation performance changes as we vary the number of slots during evaluation. We note that in the baseline and layer norm variants, segmentation performance deteriorates when they are presented with more than nine slots, while our proposed normalizations appear to generalize well to these changes. In particular, we observe that both of our proposed normalizations outperform the baseline when the autoencoders are evaluated with 11 slots, as is done in (Locatello et al., 2020).
@@
In this setting, the studied variants seem to perform more comparably than before w.r.t. foreground segmentation performance (Subfigures 2(a)-2(c)). Notwithstanding, we again note that the performance of the baseline and layer norm variants starts to suffer as we add additional slot latents during inference. In contrast, the performance of the two proposed variants remains more stable. In general, the foreground segmentation performance is lower than during training with few slots (Subfigures 1(a)-1(c)). All models trained with our proposed methods learn to segment the background into a single slot, leading to high overall segmentation performance (Subfigures 2(d)-2(f)). One model using the baseline normalization exhibits this behavior, while none of the models using layer normalization do so.
@@ 5.2 Object Discovery on MOVi-C
In our first set of experiments, we closely follow the setup detailed in (Seitzer et al., 2023) and train the autoencoders on the full MOVi-C10 dataset, using 11 slots. As in the previous subsection, we investigate how the foreground segmentation performance develops as we vary the number of slot latents during inference. For each model variant, we perform 5 training runs with different seeds. For sake of consistency with the other experiments, we evaluate the trained models both on the MOVi-C10 test set and on the filtered MOVi-C6 dataset. In Subfigures 3(a) and 3(b), we observe that our proposed normalizations generally lead to improved foreground segmentation performance compared to the baseline and layer normalization across all slot counts. In particular, both proposed normalizations outperform the baseline and layer normalization variant with 11 slots on the MOVi-C10 dataset, as can be observed in Subfigures 3(a) and 3(b). As in our previous experiments, we note that foreground segmentation performance starts to suffer as the baseline variant is provided with excess slots during inference. While our proposed methods exhibit a similar behavior in these experiments, we note that the deterioration progresses at a slower rate. Additionally, we observe in Subfigure 3(c) that performance is improved across smaller object counts.
@@
The behavior of overall segmentation performance is shown in Subfigures 3(d)-3(f). In line with our previous observations, we note that performance suffers for all variants when excess slots are present during inference. While baseline, layer normalization and batch normalization yield comparable results w.r.t. this metric, the weighted sum variant performs noticably better.
@@
We illustrate in Subfigures 4(a)-4(c) the behavior of the foreground segmentation performance. As in our previous experiments, we observe that both of our proposed normalizations outperform the baseline when evaluated on the MOVi-C10 test set with 11 slots. Interestingly, it can be noted that performance also improves over the models trained on the MOVi-C10 dataset with 11 slots (Figure 3). Again, we point out that excess slot latents during inference lead to a substantial deterioration of foreground segmentation performance in the baseline and layer norm models.
In Subfigures 4(d)-4(f), we plot the behavior of overall segmentation quality. Compared to the previous experiment, varying slot count has a lesser impact on overall segmentation performance for all methods. While the differences seem unsubstantial, the baseline appears to perform best w.r.t. this metric.
@@
In line with the experiments on the CLEVR dataset, we turn to a set of experiments that investigates the impact of excess slots during training. Similar to the corresponding setup in Subsection 5.1, we train the models on the filtered MOVi-C6 dataset, but provide them with 11 slots during training. We generally observe that both the weighted sum normalization and the batch normalization lead to improved foreground segmentations compared to the baseline and layer normalization, especially at higher slot count, as can be concluded from Subfigures 5(a) and 5(b).
@@
(b) F-ARI ($\uparrow$) on MOVi-C10, number of slots varies inference.
@@
Subfigure 5(c) additionally demonstrates once again that our proposed normalizations appear to perform at least as well as the baseline across all object counts.
In Subfigures 5(d)-5(f) we find that increased slot count during inference harms overall segmentation quality. The weighted sum variant performs best, although the variablity in ARI appears large.
@@
We now investigate how the previously described models (which were trained on MOVi-C) transfer to the MOVi-D dataset. We recall that scenes of the MOVi-D dataset may contain up to 23 objects. Hence, the MOVi-D dataset allows us to study how the trained models behave when slot- and object-count are increased significantly during inference. We evaluate the models with 24 slots. In Table 1, we show the MOVi-D zero-shot performance of all 12 model variants we trained on MOVi-C. The batch norm variant performs best w.r.t. our main metric, the F-ARI. This holds true both when comparing all 12 variants to each other, and when considering the subsets of models obtained by only considering variants with any fixed annotation $(O,K)$. Compared to their batch normalization counterparts, the weighted sum models perform similarly w.r.t. foreground ARI, performing only slightly worse. Both the weighted sum normalization and the batch normalization produce markedly better foreground segmentations than the baseline and layer norm variants. With respect to overall segmentation quality (ARI), none of the normalization variants consistently outperform the others, which is in line with our observations on MOVi-C. For any fixed normalization method, the $(6,7)$ variant performs better w.r.t. F-ARI than the $(10,11)$ and $(6,11)$ variants. This illustrates that, firstly, training on filtered training sets with few objects can improve performance during inference on scenes with many objects; secondly, avoiding excess slots during training appears to be important, which is an observation that has been made in (Locatello et al., 2020) and which we also note when comparing Figures 1 and 2. This effect underlines the importance of strong slot-count generalization capabilities. We posit that training at low object- and slot-counts reduces the number of "unoccupied" slots during training, thereby tightening the representational slot-bottleneck, which has been postulated to encourage object-centricness (Locatello et al., 2020; Stange et al., 2023).

### v2

@@ Abstract
Object-centric scene decompositions are important representations for downstream tasks in fields such as computer vision and robotics. The recently proposed Slot Attention module, already leveraged by several derivative works for image segmentation and object tracking in videos, is a deep learning component which performs unsupervised object-centric scene decomposition on input images. It is based on an attention architecture, in which latent slot vectors, which hold compressed information on objects, attend to localized perceptual features from the input image. In this paper, we demonstrate that design decisions on normalizing the aggregated values in the attention architecture have considerable impact on the capabilities of Slot Attention to generalize to a higher number of slots and objects as seen during training. We propose and investigate alternatives to the original normalization scheme which increase the generalization capabilities of Slot Attention to varying slot and object counts, resulting in performance gains on the task of unsupervised image segmentation. The newly proposed normalizations represent minimal and easy to implement modifications of the usual Slot Attention module, changing the value aggregation mechanism from a weighted mean operation to a scaled weighted sum operation.
@@ 2.1 Slot Attention
For the attention mechanism, an unnormalized $N\times K$ matrix $\bm{M}$ of dot products is formed from the keys $\bm{k}_{n}:=k(\bm{x}_{n})$ and queries $\bm{q}_{k}:=q(\bm{\theta}_{k})$. On each row of $\bm{M}$, a Softmax operator is then applied, yielding $\bm{\Gamma}=(\gamma_{n,k})\in[0,1]^{N\times K}$: $M_{n,k}:=\frac{1}{\tau}k(\bm{x}_{n})^{\top}q(\bm{\theta}_{k})=\frac{1}{\tau}\bm{k}_{n}^{\top}\bm{q}_{k}$ (1) $\gamma_{n,k}:=\frac{\exp{M_{n,k}}}{\sum_{k^{\prime}=1}^{K}\exp{M_{n,k^{\prime}}}}.$ (2) With this, each row $\bm{\gamma}_{n,:}$ may be interpreted as the probability of an input token $n$ to be assigned to a particular slot $k$. The constant $\tau$ corresponds to a temperature parameter which is chosen to be $\sqrt{D}$.
A linear map $v:\mathbb{R}^{D_{\text{input}}}\to\mathbb{R}^{D}$ extracts values from the input tokens and the matrix $\bm{\Gamma}$ is used to accumulate values into unnormalized slot-wise update codes:
@@ 3 Slot Attention and von Mises-Fisher Mixture Model Parameter Estimation
Many works (Locatello et al., 2020; Chang et al., 2022b; a; Kirilenko et al., 2023) compare Slot Attention to expectation maximization (Dempster et al., 1977; Bishop, 2006) (EM) in Gaussian mixture models, i.e. to soft k-means clustering. We, however, connect it with expectation maximization in a mixture model of von Mises-Fisher (vMF) distributions (Banerjee et al., 2003), since Slot Attention uses a bilinear form on slots and inputs as a scoring function instead of the negative Euclidean distance. In this section, we make the parallel between Slot Attention and EM explicit by performing EM parameter estimation in a vMF mixture model and relating each step to the corresponding step in Slot Attention. We will then view the weighted mean, layer norm, and weighted sum normalization variants in the context of this analogy and compare them.
@@ 3.2 Comparing Normalizations in EM Analogy
In the weighted mean case, the aggregated values $\bm{u}_{k}$ can hold sufficient information to extract the quantities in (7) and (8) if $D={D_{\text{input}}}$ holds. Assuming that the value map is the identity, the right-hand side of (7) may be computed as $\bm{u}_{k}/\|\bm{u}_{k}\|_{2}$. However, it is not clear how $\pi_{k}$ could be computed from $\bm{u}_{k}$. Indeed, we show in Proposition 1 that there can be no general formula as for the weighted sum case that generalizes without exception across slot-counts. We provide a proof in Appendix C by constructing some explicit slot settings which demonstrate that a hypothetical function $f$ can not map every update code to a corresponding unique scalar.
@@
for arbitrary $K\geq 1$, arbitrary slots $\tilde{\bm{\theta}}_{1},...,\tilde{\bm{\theta}}_{K}$ and resulting normalized update codes $\bm{u}_{1},...,\bm{u}_{k}$.
@@
Since the assumptions of Proposition 2 only exclude a parameter subset of zero volume, its conclusion likely holds in practice during training.
Hence, the weighted sum variant may also be seen as a generalization of the weighted mean normalization: the update codes from weighted sum normalization hold sufficient information such that weighted-mean-normalized update codes can be recovered from them (e.g. by the update network if it has sufficient capacity). The reverse is not true, as demonstrated in Proposition 1.
@@ 4.1 Weighted Sum Normalization with Fixed Scaling
Here, we first use the triangle inequality, followed by the crude estimate $|v_{n,d}|\leq\max_{n}|v_{n,d}|$ for all $n$. Finally, we used the fact that $\sum_{n=1}^{N}\gamma_{n,k}\leq N$ holds, since $\bm{\Gamma}$ is row-stochastic with $N$ rows. While this provides an argument for our choice of $C$ which is a suitable heuristics across tasks, we hypothesize that task-specific tuning of this hyperparameter may be beneficial.
@@ 4.2 Weighted Sum Normalization with Batch Scaling
where $\epsilon>0$ is a small constant. We stress that the values $m$ and $v$ are computed for each mini-batch in the first Slot Attention iteration and are therefore independent of $j$. Moreover, gradients flow through $m$ and $v$. We cache an exponential moving average of the batch statistics during training and use it during inference. Hence, during inference, the normalization in equation (13) is simply an affine transformation with fixed weights, thereby closely resembling the weighted sum normalization presented in the previous subsection. While the weighted sum normalization is linear and not affine, we note that this distinction does not impact the capacity of the model. Indeed, the vectors $\bm{u}_{k}$ are also affinely transformed within the update network, therefore making any previous affine or linear transformation redundant from a capacity perspective. Notwithstanding, the normalizations presented here will significantly alter the trainig trajectory of the models. Batch normalization, in particular, has previously been shown to remedy the problem of saturating activations and vanishing gradients (Ioffe & Szegedy, 2015; Pascanu et al., 2013), which may arise from improper normalization (Glorot & Bengio, 2010). We provide pseudocode for the two proposed normalization variants in Appendix G.
@@ 5 Experiments
We investigate the proposed normalizations on unsupervised object discovery tasks. To this end, we train autoencoders on the CLEVR (Johnson et al., 2017) and MOVi-C (Greff et al., 2022) datasets, utilizing autoencoder architectures that have been described in (Locatello et al., 2020) and (Seitzer et al., 2023), respectively. Additional results for a property prediction task can be found in Appendix I. We provide visualizations of scene segmentations in Appendix H. We will give a brief overview on our experimental setup in the following and refer to the supplementary material for more details [^1] Code is available at https://github.com/EmbodiedVision/slot_attention_normalization..
@@ 5.1 Object Discovery on CLEVR
In this first set of experiments, we follow (Locatello et al., 2020) as closely as possible and train the previously described CNN-based architecture on the CLEVR6 dataset with 7 slots. We compare the baseline and layer normalizations proposed in (Locatello et al., 2020) to the methods discussed in Section 4. For each variant, we perform 5 training runs with different seeds. The trained models are evaluated on the CLEVR6 and CLEVR10 test sets.
The baseline and layer norm variants lead to object-centric behavior for all 5 seeds. For the weighted sum normalization and the batch norm variant, however, we encounter two runs each in which the autoencoders decompose the input spatially instead of object-wise. Following (Locatello et al., 2020), we omit these runs in our analysis.
@@
In Subfigures 1(a) and 1(b), we illustrate how the foreground segmentation performance changes as we vary the number of slots during evaluation. We note that in the baseline and layer norm variants, segmentation performance deteriorates when they are presented with more than nine slots, while our proposed normalizations appear to generalize well to these changes. In particular, we observe that both of our proposed normalizations outperform the baseline when the autoencoders are evaluated with 11 slots, as is done in (Locatello et al., 2020). We show qualitative results of the different model variants at a high slot count in Figure 2 and refer to Appendix H for more visualizations.
Input Baseline Layer Norm Weighted Sum Batch Norm
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Qualitative results on a CLEVR10 images, showing reconstructions and (soft) segmentations. The models are trained on CLEVR6 with 7 slots and evaluated with 21 slots.
@@
In this setting, the studied variants seem to perform more comparably than before w.r.t. foreground segmentation performance (Subfigures 3(a)-3(c)). Notwithstanding, we again note that the performance of the baseline and layer norm variants starts to suffer as we add additional slot latents during inference. In contrast, the performance of the two proposed variants remains more stable. In general, the foreground segmentation performance is lower than during training with few slots (Subfigures 1(a)-1(c)). All models trained with our proposed methods learn to segment the background into a single slot, leading to high overall segmentation performance (Subfigures 3(d)-3(f)). One model using the baseline normalization exhibits this behavior, while none of the models using layer normalization do so.
@@ 5.2 Object Discovery on MOVi-C
In our first set of experiments, we closely follow the setup detailed in (Seitzer et al., 2023) and train the autoencoders on the full MOVi-C10 dataset, using 11 slots. As in the previous subsection, we investigate how the foreground segmentation performance develops as we vary the number of slot latents during inference. For each model variant, we perform 5 training runs with different seeds. For sake of consistency with the other experiments, we evaluate the trained models both on the MOVi-C10 test set and on the filtered MOVi-C6 dataset.
In Subfigures 4(a) and 4(b), we observe that our proposed normalizations generally lead to improved foreground segmentation performance compared to the baseline and layer normalization across all slot counts. In particular, both proposed normalizations outperform the baseline and layer normalization variant with 11 slots on the MOVi-C10 dataset, as can be observed in Subfigures 4(a) and 4(b). As in our previous experiments, we note that foreground segmentation performance starts to suffer as the baseline variant is provided with excess slots during inference. While our proposed methods exhibit a similar behavior in these experiments, we note that the deterioration progresses at a slower rate. Additionally, we observe in Subfigure 4(c) that performance is improved across smaller object counts.
@@
The behavior of overall segmentation performance is shown in Subfigures 4(d)-4(f). In line with our previous observations, we note that performance suffers for all variants when excess slots are present during inference. While baseline, layer normalization and batch normalization yield comparable results w.r.t. this metric, the weighted sum variant performs noticably better.
@@
We illustrate in Subfigures 5(a)-5(c) the behavior of the foreground segmentation performance. As in our previous experiments, we observe that both of our proposed normalizations outperform the baseline when evaluated on the MOVi-C10 test set with 11 slots. Interestingly, it can be noted that performance also improves over the models trained on the MOVi-C10 dataset with 11 slots (Figure 4). Again, we point out that excess slot latents during inference lead to a substantial deterioration of foreground segmentation performance in the baseline and layer norm models.
In Subfigures 5(d)-5(f), we plot the behavior of overall segmentation quality. Compared to the previous experiment, varying slot count has a lesser impact on overall segmentation performance for all methods. While the differences seem unsubstantial, the baseline appears to perform best w.r.t. this metric.
@@
In line with the experiments on the CLEVR dataset, we turn to a set of experiments that investigates the impact of excess slots during training. Similar to the corresponding setup in Subsection 5.1, we train the models on the filtered MOVi-C6 dataset, but provide them with 11 slots during training. We generally observe that both the weighted sum normalization and the batch normalization lead to improved foreground segmentations compared to the baseline and layer normalization, especially at higher slot count, as can be concluded from Subfigures 6(a) and 6(b).
@@
Subfigure 6(c) additionally demonstrates once again that our proposed normalizations appear to perform at least as well as the baseline across all object counts. In Subfigures 6(d)-6(f) we find that increased slot count during inference harms overall segmentation quality. The weighted sum variant performs best, although the variablity in ARI appears large.
@@
We now investigate how the previously described models (which were trained on MOVi-C) transfer to the MOVi-D dataset. We recall that scenes of the MOVi-D dataset may contain up to 23 objects. Hence, the MOVi-D dataset allows us to study how the trained models behave when slot- and object-count are increased significantly during inference. We evaluate the models with 24 slots. In Table 1, we show the MOVi-D zero-shot performance of all 12 model variants we trained on MOVi-C. The batch norm variant performs best w.r.t. our main metric, the F-ARI. This holds true both when comparing all 12 variants to each other, and when considering the subsets of models obtained by only considering variants with any fixed annotation $(O,K)$. Compared to their batch normalization counterparts, the weighted sum models perform similarly w.r.t. foreground ARI, performing only slightly worse. Both the weighted sum normalization and the batch normalization produce markedly better foreground segmentations than the baseline and layer norm variants. With respect to overall segmentation quality (ARI), none of the normalization variants consistently outperform the others, which is in line with our observations on MOVi-C.
For any fixed normalization method, the $(6,7)$ variant performs better w.r.t. F-ARI than the $(10,11)$ and $(6,11)$ variants. This illustrates that, firstly, training on filtered training sets with few objects can improve performance during inference on scenes with many objects; secondly, avoiding excess slots during training appears to be important, which is an observation that has been made in (Locatello et al., 2020) and which we also note when comparing Figures 1 and 3. This effect underlines the importance of strong slot-count generalization capabilities. We posit that training at low object- and slot-counts reduces the number of "unoccupied" slots during training, thereby tightening the representational slot-bottleneck, which has been postulated to encourage object-centricness (Locatello et al., 2020; Stange et al., 2023).